1. When 10g of N2 and 7g H2 are allowed to react. Determine the limiting reactant and excess reactant in this reaction.


N_{2}+H_{2} ->NH_{3}


Given:

Find:

Solution:


1
Expert's answer
2022-04-11T10:08:04-0400

Given:

Mass of N2 = 10 g

Mass of H2 = 7 g


Find:

The limiting reactant is ???

The excess reactant is ???


Solution:

The molar mass of N2 is 28.02 g/mol

The molar mass of H2 is 2.016 g/mol


Calculate the moles of each reactant:

Moles of N2 = (10 g N2) × (1 mol N2 / 28.02 g N2) = 0.357 mol N2

Moles of H2 = (7 g H2) × (1 mol H2 / 2.016 g H2) = 3.472 mol H2


Balanced chemical equation:

N2 + 3H2 → 2NH3

According to the chemical equation above:

1 mol of N2 reacts with 3 mol of H2

Thus, 0.357 mol of N2 reacts with:

(0.357 mol N2) × (3 mol H2 / 1 mol N2) = 1.071 mol H2

However, initially there is 3.472 of H2 (according to the task)

Therefore, N2 acts as limiting reactant and H2 is excess reactant


Answer:

The limiting reactant is N2

The excess reactant is H2